OpenStudy (anonymous):

Solve: 2(x - 3) + 2(x - 7) = 40 A. x = 12.5 B. x = 15 C. x = 5 D. no solution E. infinitely many solutions

sam (.sam.):

2(x - 3) + 2(x - 7) = 40 distribute 2x-6+2x-14=40 4x-20=40 solve